# A note on cyclic non-MDS matrices

### v1

@@ 6 Results on orthogonal cyclic matrices
We begin this section with an alternative proof of theorem 5.4 using the fact that the permutation matrices are orthogonal. This proof holds for a more general class, i.e., for cyclic matrices. Although this proof is compact, it does not describe any properties of cyclic matrices like the determinant or scaler structure. But for the sake of completeness, we record the proof here.
@@ 7 Conclusion
In this article, we focus on the structure of cyclic matrices with an application of it in the construction of MDS matrices. We present some results on orthogonal $g$-circulant matrices with MDS property. Additionally, we provide examples of cyclic MDS matrices with orthogonal property. Even though, a general structure of cyclic matrices is provided, a study of cyclic matrices of orders other than $2^{d}\times 2^{d}$ with orthogonal property is still an interesting problem.

### v2

@@ 6 Results on orthogonal cyclic matrices
We begin this section with an alternative proof of theorem 5.4 using the fact that the permutation matrices are orthogonal. This proof holds for a more general class, i.e., for cyclic matrices. Although this proof is compact, it does not describe any properties of cyclic matrices like the determinant or scalar structure. But for the sake of completeness, we record the proof here.
@@
Type of MDS Matrix Order Existence Reference
circulant, orthogonal $2^{d}\times 2^{d}$ do not exist [20]
circulant, orthogonal other orders may exist [20]
left- circulant, orthogonal $2^{d}\times 2^{d}$ do not exist [23]
g-circulant, orthogonal $2^{d}\times 2^{d}$ do not exist this paper
cyclic, orthogonal $2^{d}\times 2^{d}$ do not exist this paper
cyclic, orthogonal other orders exist this paper
Table 1: List of pre-existing results and new results over finite fields of characteristic $2$.
@@ 7 Conclusion
In this article, we focus on the structure of cyclic matrices and their application in the construction of MDS matrices. Cyclic matrices are closely related to circulant matrices and therefore offer flexible and efficient hardware implementations. This makes the study of cyclic matrices both practically relevant and theoretically significant. Due to these foundational properties, we study the characteristics of cyclic matrices and present new results on orthogonal $g$-circulant matrices with the MDS property. Additionally, we provide explicit examples of cyclic MDS matrices that also possess the orthogonal property.
While this article establishes a general structure for cyclic matrices and cyclic orthogonal MDS matrices, it primarily focuses on matrices of size $2^{d}\times 2^{d},$ where $d$ is a positive integer. However, the study of cyclic matrices with orders other than powers of two remains an open and intriguing area of research. Additionally, investigating the generalization of orthogonal properties, as well as involutory and semi-involutory properties in cyclic matrices, presents another promising direction for future study.